OpenStudy (anonymous):

Give the measure of the angle in radians and degrees arcsin (-square root of 2/2

OpenStudy (anonymous):

@satellite73 please help

OpenStudy (anonymous):

@ganeshie8 please help

OpenStudy (anonymous):

@tkhunny please help

OpenStudy (tkhunny):

Do you have a "Unit Circle" to look at? You should be quite familiar with \(-\pi/4\). Why negative and not positive?

OpenStudy (anonymous):

Where did you get -pi/4

OpenStudy (anonymous):

Why did you use pi/4 and not something like 3pi/4?

OpenStudy (tkhunny):

Did you find a "Unit Circle", yet?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

For one, the arcsin function typically is defined only for \(-\pi/2\le arcsin(x)\le\pi/2\). This rules out quite a few options.

OpenStudy (tkhunny):

What does your unit circle tell you about the sine of \(-\pi/4\)?

OpenStudy (anonymous):

Its 45 degrees and is (square root of 2/2, square root of 2/2

OpenStudy (tkhunny):

No, that's \(\pi/4\). We need \(-\pi/4\) When you find it, it will say, \(\sin(-45º) = -\sqrt{2}/2\). This should look quite tantalizing in light of the original problem statement.

OpenStudy (anonymous):

Would that be 7pi/4?

OpenStudy (tkhunny):

Well, that is the same place, but it is NOT the answer to the original problem statement. This is correct: \(sin(-\pi/4) = -\sqrt{2}/2\) \(sin(7\pi/4) = -\sqrt{2}/2\) However, and this may come as a surprise: \(arcsin(-\sqrt{2}/2) = -\pi/4\) and is absolutely NOT \(7\pi/4\). It is a Range Issue. The arcsin function does not contain \(7\pi/4\) in its Range.

OpenStudy (anonymous):

Ok I understand, if the range was from -infinity to infinity, 7pi/4, would be a valid answer, but since it isn't you just have to slap on the negative to pi/4

OpenStudy (tkhunny):

Not quite, but that may get you to the next problem.

OpenStudy (anonymous):

thanks for helping